Minutes of July 8, 2009 meeting

Andy Martinez (Coordinator), Elsa Alves, Kimberly Kaiser, Rob Krack, Bill Lynch, Shirley Peck, Eddie Suarez (Recorder), Jan Reinhart
Teri McNally, Penny Weniger, Mary Brodman, Eugene McElroy, Brian Stubbs, Melody Tomaszewicz,


Information Sharing

Andy (Coordinator): Andy is working on the quarterly missing book lists. Two lists will be released: the final search for items charged to missing before July 1, 2008 and the search for items charged to missing from July 1, 2008 to July 15, 2009 report.
An inventory of the classed collections at the Physics Library should be completed soon. Once that inventory is completed, one will start at the Chemistry Library.

Elsa (Dana): Elsa reported that the Stacks Collection shift is ¾ completed. The shift has slowed due to lack of student workers.

Shirley (Alex): Shirley reported on a problem she experienced with in-transit items during the processing of "in-process" items.

Rob (Alexander): Rob reported on the preparation for the Reference collection shift. 140 rent-a-carts were rented for the reference collection, which was stored in basement while carpeting was being installed. Ref folio (35 cm) and E Folio (45 cm) size classification were instituted in a new shelving location. Rob reported an upswing in shelving due to summer classes. Mini-shifts of the Stacks collection are taking place on 2B and 3rd floors. Finally,the in-house statistical pages for Alexander's collection management were completed.

Rob trained students to shelve in Government Documents. A shift is taking place in the Doc Law collection. 80 moldy books were removed from Gov Docs as a result of high humidity in the area.

Rob also reported that Erin Uhl is the new associate in preservation and that CM members should report any preservation issues to her.

Kim (Kilmer): Kim reported that the Reference Collection was weeded and the smaller collection was shifted. The smaller Reference Collection has opened up space for users of the libraries.

Bill (Douglass): Bill reported on a weeding project in the East Room (Stacks Collection) to prepare for the collection shift in the West room starting in September.

Eddie (LSM): Eddie was able to shelve all of the year-end returns in the Stacks and Periodical Collections. He also received a weeding list from librarians. Eddie continues a mini-shift of the Periodical Collection.

Jan (Media): Jan reported on working with Jane Ott on breaking up DVD volume sets in Media into individual records.

Water Emergency Powerpoint (Brief Recap)

Rob, Kim and Eddie have completed the third draft of the PowerPoint. They discussed a plan for new updated pictures and new information to be added to the slides. The group is investigating how to add voice tracks to the slides, with the presentation meant to be viewed online. They might coordinate a live presentation of how to apply live plastic to a library collection that has been impacted by water.

Next Meeting

August/2009 meeting is cancelled

Wednesday, 9/9/2009
2:30 - 4:30pm
Location TBA

URL: http://www.libraries.rutgers.edu/rul/staff/groups/collection_mgt_group/minutes/collection_mgt_group_09_07_08.shtml
Website Feedback  |  Privacy Policy
© Copyright 1997-, Rutgers University Libraries